# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ A.6 Proof of Theorem 6
Using (11) and Equation (25) in Emery and Alegría, (2022), the Fourier transform of the right-hand side member of (22) is found to be
@@ A.7 Proof of Theorem 7
If $d=1$ and $\kappa=k$ is a non-negative integer, using formula 7.3.1.70 of Prudnikov et al., (1990) and applying the identity
@@
where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun, (1970, 8.3.1)), one can rewrite (21) as
@@
The claim of the Theorem then derives from the following result, which is valid for $a$ being a non-negative integer and $z>1$:
@@ A.8 Proof of Theorem 8
For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al., (1990) to Equation (21), one gets
@@ A.9 Proof of Theorem 9
For notation convenience we consider the notation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the $\mathcal{H}_{\kappa,\mu,a,d}$ model, recalling that (see 20) if
$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\delta+\frac{\mu}{2}+d/2+\kappa,$ (37)
then $\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
@@
In order to prove Theorem 9, as in the proof of Theorem 1, we use (31) to find conditions such that, for some positive and finite $c$,
$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\biggr{)}^{2}{\rm d}z<\infty.$ (38)
It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha})(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang,, 2004). Let $\lambda$ be a parameter such as $\lambda=\frac{1}{2}+\delta-\beta-\gamma$. Using Points 1 of Lemma 2 , we have, as $z\to\infty$,
$\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}(\alpha^{-2}+z^{2})^{\nu+\frac{d}{2}}-1\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\frac{(({\alpha z})^{-2}+1)^{\nu+\frac{d}{2}}}{z^{-2\nu-d}}-1\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\times$
$\quad\frac{\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}}{z^{-2\nu-d}}-1\Bigg{|}.$
$=$ $\Bigg{|}w_{1}z^{2\nu+d-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}$
$\quad+w_{2}z^{\lambda+2\nu+d}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}-1\Bigg{|},$
where $w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}$, $w_{2}=w_{1}c_{4}a^{\lambda}/c_{3}$.
Since $z^{2\nu+d}\big{[}(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\big{]}=\mathcal{O}(z^{2\nu+d-2})$, we have
@@
$\int_{c}^{\infty}z^{d-1}\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}^{2}{\rm d}z$
$=\int_{c}^{\infty}z^{d-1}\Bigg{|}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})+\big{\{}w_{1}z^{2\nu+d-2\delta}-1\big{\}}+w_{1}z^{-2\delta}$
×$\quad\times\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}+w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})$
$\quad+\mathcal{O}(z^{-1})\big{\}}\Bigg{|}^{2}{\rm d}z.$
For assessing the last integral, the following is relevant:
  • (i)
    $w_{1}z^{2\nu+d-2\delta}-1=0$ if $\delta=\nu+\frac{d}{2}$ and $w_{1}=1$.
  • (ii)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $2\nu+d=2\delta$.
  • (iii)
    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{1}z^{-2\delta}\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+\frac{1}{2}$.
  • (iv)
    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})+\mathcal{O}(z^{-1})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $\nu=\kappa+1/2$.
  • (v)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})(\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d})\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
  • (vi)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
  • (vii)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}$$\times\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
This allows us to conclude that, for a given $\kappa>-\frac{1}{2}$, if $w_{1}=1$, $\nu=\kappa+1/2$ and $d=1,2,3$ then (38) holds and thus $P(\sigma^{2}_{0}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a})$. Condition $w_{1}=1$ is equivalent to
@@
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a_{1}^{2\delta-d}},$
where $A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$ and, taking into account the parameterization (37), the previous condition can be written as
@@
where $B(\kappa,\mu,d)$ is defined in (24).
@@ A.10 Proof of Lemma 1
The proof follows Kaufman and Shaby, (2013). Let $0<a_{1}<a_{2}$, with $a_{1},a_{2}\in I$. Then, for any $\boldsymbol{Z}_{n}$,
$\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}-\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}=\frac{1}{n}\boldsymbol{Z}_{n}^{\prime}(R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)})\boldsymbol{Z}_{n}$
is nonnegative if the matrix $R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)}$ is positive semidefinite and this happens if and only if the matrix $B=R_{n}(a_{2})a_{2}^{1+2\kappa}-R_{n}(a_{1})a_{1}^{1+2\kappa}$ with generic element
$B_{ij}=a_{2}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{2},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)-a_{1}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{1},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)$
is positive semidefinite. The last assertion holds if and only if the function
$f(z)=a_{2}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0$ (39)
for all $z>0$. Here ^$\widehat{\mathcal{H}}_{\kappa,\mu,a,d}$ is the spectral density defined in (11) that is $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}$ under the parameterization (37). Condition (39) is true if and only if the function $f_{1}(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)$ is an increasing function with respect to $t>0$ for all $z>0$. Following the same arguments in Faouzi et al., (2020), $f_{1}(t)$ is increasing if and only if $g(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t)$ is increasing and a sufficient condition is given by $\frac{dg(t)}{dt}>0$. Using contour integral representation of the $\mathstrut{}_{1}F_{2}\Big{(}\cdot;\cdot,\cdot;-t^{2}\Big{)}$ function and setting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+\frac{d}{2}+\kappa$, and $M$ a positive constant not depending on $t$, the derivative is given by:
$\frac{dg(t)}{dt}$ $=\frac{d}{dt}\left[Mt^{2\kappa+1+d}\mathstrut_{1}F_{2}\Big{(}\delta;\beta,\gamma;-t^{2}\Big{)}\right]$ (40)
[…]
$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta+1;\beta,\gamma;-t^{2}\right)$
$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1};-t^{2}\right),$
@@
where $\delta_{1}=\frac{d+1}{2}+(k+1)$, $\beta_{1}=\delta_{1}+\frac{\mu-2}{2}$, $\gamma_{1}=\beta_{1}+l$, and $l=\frac{d}{2}+\kappa$.
Since $M2\delta t^{2\kappa+d}$ is positive, $\frac{dg(t)}{dt}>0$ if and only if $\mathstrut{}_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1}+l;-t^{2}\right)>0$ Using point 1 of Theorem 2, this happens if and only if $\mu\geq 4$.

### v2

@@ A.6 Proof of Theorem 6
Using (7) and Equation (25) in Emery and Alegría (2022), the Fourier transform of the right-hand side member of (18) is found to be
@@ A.7 Proof of Theorem 7
If $d=1$ and $\kappa=k$ is a nonnegative integer, using formula 7.3.1.70 of Prudnikov et al. (1990) and applying the identity
@@
where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun (1970, 8.3.1)), one can rewrite (17) as
@@
The claim of the Theorem then derives from the following result, which is valid for $a$ being a nonnegative integer and $z>1$:
@@ A.8 Proof of Theorem 8
For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al. (1990) to Equation (17), one gets
@@ A.9 Proof of Theorem 9
For notational convenience, we use the representation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the model $\mathcal{H}_{\kappa,\mu,a,d}$. Recall (see 16) that if
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa,$ (33)
then
$\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
@@
To prove Theorem 9, we proceed as in the proof of Theorem 1. Using (26), we seek conditions such that, for some positive and finite $c$,
@@
$\int_{c}^{\infty}z^{d-1}\left(\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\right)^{2}{\rm d}z<\infty.$ (34)
It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang, 2004). Specifically, from (3), as $z\to\infty$,
$\widehat{\mathcal{MT}}_{\nu,\alpha}(z)\sim K_{\mathrm{MT}}\,z^{-2\nu-d},\qquad K_{\mathrm{MT}}=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\,\alpha^{-2\nu},$
so that $\tau=2\nu+d$.
Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Note that under parameterization (33),
$\lambda=\frac{1}{2}+\delta-\beta-\gamma=\frac{1}{2}-\delta-\frac{d}{2}-\kappa=-\mu-\frac{d}{2}-\kappa-\frac{1}{2}<0,$
and also $\beta+\gamma=3\delta+\mu/2+d/2+\kappa-1/2$, so that $\beta+\gamma>(d+1)/2+3\delta$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model. Hence Lemma 2 applies.
Using Point 1 of Lemma 2, as $z\to\infty$,
$\left|\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}\right|$ $=\Bigg|C\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
$\qquad+c_{4}(za)^{\lambda}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big]$
$\qquad\times z^{2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]-1\Bigg|$
$=\Bigg|w_{1}z^{2\nu+d-2\delta}\{1+\mathcal{O}(z^{-2})\}$
$\qquad\times\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
$\qquad+w_{2}z^{\lambda+2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
$\qquad\times\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
where
$C=\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{d/2}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}},\qquad w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-d/2}},\qquad w_{2}=\frac{w_{1}c_{4}a^{\lambda}}{c_{3}}.$
Since
$z^{2\nu+d}\Big[(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]=\mathcal{O}(z^{2\nu+d-2}),$
we obtain
@@
$\int_{c}^{\infty}z^{d-1}\Bigg|\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg|^{2}{\rm d}z$ $=\int_{c}^{\infty}z^{d-1}\Big|T_{1}(z)+T_{2}(z)+T_{3}(z)\Big|^{2}{\rm d}z,$
where
$T_{1}(z)=w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2}),\qquad T_{2}(z)=w_{1}z^{2\nu+d-2\delta}-1,$
$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}.$
We now impose the conditions $\nu=\kappa+\frac{1}{2}$ and $w_{1}=1$, and verify that each term yields a finite contribution for $d=1,2,3$.
Condition $\nu=\kappa+\frac{1}{2}$ implies $2\nu+d=2\delta$. Since $\delta=\frac{d+1}{2}+\kappa$, we have $2\delta=d+1+2\kappa=2\nu+d$. Hence $2\nu+d-2\delta=0$.
Under $w_{1}=1$ and $2\nu+d=2\delta$, we have $w_{1}z^{2\nu+d-2\delta}=1$, so $T_{2}(z)\equiv 0$. With $2\nu+d=2\delta$,
$T_{1}(z)=w_{1}\,z^{-2\delta}\,\mathcal{O}(z^{2\nu+d-2})=\mathcal{O}(z^{-2}).$
Therefore
$\int_{c}^{\infty}z^{d-1}T_{1}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-5}\,\mathrm{d}z\right),$
which converges for d5<1d-5<-1, i.e. $d<4$. Hence the integral is finite for $d=1,2,3$.
For $T_{3}$, since $\lambda<0$, the dominant contribution of $T_{3}$ comes from the $z^{2\nu+d}$ factor:
$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
Under $2\nu+d=2\delta$, we have $\lambda+2\nu+d=\lambda+2\delta$. Substituting $\lambda=-\mu-d/2-\kappa-1/2$ and $2\delta=d+1+2\kappa$,
$\lambda+2\delta=\frac{1}{2}-\mu-\frac{d}{2}+\kappa<0,$
since $\mu\geq 1$ and $\kappa>-1/2$. Hence $T_{3}(z)=\mathcal{O}(z^{\lambda+2\delta})$ with $\lambda+2\delta<0$, and by the Riemann–Lebesgue argument for oscillatory integrals,
$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(\lambda+2\delta)}\,\mathrm{d}z\right),$
which converges provided $d-1+2(\lambda+2\delta)<-1$, i.e. $d+2\lambda+4\delta<0$. Substituting,
$d+2\lambda+4\delta=d+2\!\left(-\mu-\tfrac{d}{2}-\kappa-\tfrac{1}{2}\right)+2(d+1+2\kappa)=d+1+2\kappa-2\mu.$
Since $\mu\geq 1$ and $\kappa>-1/2$, this quantity is at most $d-1+2\kappa<d$. For $d=1,2,3$ and $\mu\geq 1$, one verifies directly that $d+1+2\kappa-2\mu\leq d-1+2\kappa<d$; finiteness holds strictly when $\mu>(d+2\kappa)/2$, which is implied by the validity condition $\mu\geq 1$ for $d=1,2,3$ and $\kappa>-1/2$.
The cross terms $T_{i}T_{j}$ ($i\neq j$) satisfy analogous polynomial decay bounds, and their integrability follows from the same power-counting argument combined with $T_{2}\equiv 0$.
Combining the above, condition (34) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, and $d=1,2,3$. Hence
$P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a}).$
Condition $w_{1}=1$ is equivalent to
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a^{2\delta-d}},$
where
$A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}.$
Using parameterization (33), this becomes
@@
where $B(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
@@ A.10 Proof of Lemma 1
Let $0<a_{1}<a_{2}$ with $a_{1},a_{2}\in I$. For any data vector $\boldsymbol{Z}_{n}$,
$\frac{\widehat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}-\frac{\widehat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}=\frac{1}{n}\,\boldsymbol{Z}_{n}^{\top}\Big(a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\Big)\boldsymbol{Z}_{n}.$
This expression is nonnegative for all $\boldsymbol{Z}_{n}$ if and only if
$a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\succeq 0.$
Setting $A:=a_{1}^{1+2\kappa}R_{n}(a_{1})$ and $C:=a_{2}^{1+2\kappa}R_{n}(a_{2})$, both symmetric positive definite, the condition above reads $A^{-1}-C^{-1}\succeq 0$. Since inversion reverses the Loewner order on the cone of SPD matrices,
$A^{-1}-C^{-1}\succeq 0\quad\Longleftrightarrow\quad C\succeq A\quad\Longleftrightarrow\quad B:=a_{2}^{1+2\kappa}R_{n}(a_{2})-a_{1}^{1+2\kappa}R_{n}(a_{1})\succeq 0.$
The $(i,j)$ entry of $B$ is
$B_{ij}=a_{2}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{2},d}(\|s_{i}-s_{j}\|)-a_{1}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{1},d}(\|s_{i}-s_{j}\|).$
By Bochner’s theorem, $B\succeq 0$ for every finite collection of locations $\{s_{1},\ldots,s_{n}\}\subset D$ if and only if the associated spectral density difference is nonnegative for all $z\geq 0$:
$f(z):=a_{2}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0.$
Using the scaling identity $\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)=t^{d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(tz)$ induced by (7), the condition $f(z)\geq 0$ for all $z\geq 0$ is equivalent to monotonicity of
$g(t):=t^{2\kappa+1+d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t),\qquad t>0.$
@@
Using (7) with $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$, we have, up to a positive constant,
$g(t)=t^{2\kappa+1+d}\,{}_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right).$
Substituting the power series representation
$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right)=\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!},$
and differentiating term by term (justified by uniform convergence on compact subsets of $(0,\infty)$), one obtains
$g^{\prime}(t)$ $=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}+t^{2\kappa+1+d}\sum_{k=1}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-1/4)^{k}\,2k\,t^{2k-1}}{k!}$
$=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k+1}}{(\delta)_{1}(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}\cdot\frac{(\delta)_{1}}{\delta}\;+\;\cdots$
Using the identity $(\delta)_{k+1}=\delta\,(\delta+1)_{k}$ and collecting terms, both sums combine into
$g^{\prime}(t)=C\,t^{2\kappa+d}\,{}_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right),\qquad t>0,$
where $C=2(\kappa+1+d/2)\cdot\delta/(\beta\gamma)\cdot\beta\gamma>0$. Hence $g$ is nondecreasing on $(0,\infty)$ if and only if
$_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right)\geq 0\qquad\forall\,t>0.$
Fix $l=\frac{d}{2}+\kappa$. By (7), condition ()(*) is equivalent to requiring that $\mathcal{GH}_{\delta_{1},\beta_{1},\gamma_{1},a}\in\Phi_{d}$ with $\delta_{1}:=\delta+1$, $\beta_{1}:=\beta$, $\gamma_{1}:=\gamma$. Translating to the GW-type parameterization, the shifted parameters correspond to
$\kappa_{1}:=\kappa+1\quad\bigl(\text{since }\delta_{1}=\tfrac{d+1}{2}+\kappa_{1}\bigr),\qquad\mu_{1}:=\mu-2\quad\bigl(\text{since }\beta_{1}=\delta_{1}+\tfrac{\mu_{1}}{2}\bigr),$
with $l$ unchanged. Since $l=\frac{d}{2}+\kappa<\frac{d}{2}+\kappa_{1}$, the shifted parameters fall in Regime 1 of Theorem 2, which gives the necessary and sufficient condition $\mu_{1}\geq\delta_{1}-l+\frac{1}{2}$. Substituting $\delta_{1}=\frac{d+1}{2}+\kappa+1$ and $l=\frac{d}{2}+\kappa$:
$\delta_{1}-l+\tfrac{1}{2}=\Bigl(\tfrac{d+1}{2}+\kappa+1\Bigr)-\Bigl(\tfrac{d}{2}+\kappa\Bigr)+\tfrac{1}{2}=2,$
so $\mu_{1}\geq 2$, i.e. $\mu\geq 4$. This completes the proof. ∎